OpenStudy (anonymous):

HELP !!! can some one tell me the interior angle and arc length of a pizzas Slices ? this is for an 16 inch pizza and there are 8 slices

OpenStudy (anonymous):

if i have a 16 inch diameter pizza cut into 8 slices what would the interior angle and arc length be of the slices?

OpenStudy (anonymous):

Circumference = pi*d Arc length = C/8 Angle = 360/8

OpenStudy (anonymous):

Ohhhh ight thnk u ! <3